TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2006-12-21 - 19:11

SCT_UA 67- UJ 67, Dec. 21st 2006

PDF RAT_20061221.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x12x21x19x03/RAT_20061221.pdf

Open Issues:

  • 21.12 EE Fan rotation to be verified
  • 21.12 Aluminum panel for the cables ducts
  • 21.12 HV Insulation test of the DQS poles
  • 18.12 ERD rack powering
  • 18.12 EOD rack powering
  • 18.12 PC powering
  • 18.12 Elettas calibration
  • 18.12 Ventilation station
  • 18.12 Cooling station test by Roese

-- Main.BorisBellesia - 21 Dec 2006 - 19:10

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-20 - 12:01

RAT Preparation Cool Down sect 78 - Dec 20th

PDF RAT_20061220.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x11x10/RAT_20061220.pdf

Next RAT meeting Tuesday, January 9th 8:30 @ P8 2889-R-009

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 20 Dec 2006 - 11:11

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2006-12-20 - 10:57

The Flags are Up

Polarity Check done in RR57-UA56

JPEG P1000172.JPG:
P1000172.JPG

JPEG P1000173.JPG:
P1000173.JPG

JPEG P1000174.JPG:
P1000174.JPG

JPEG P1000157.JPG:
P1000157.JPG

JPEG P1000154.JPG:
P1000154.JPG

JPEG P1000155.JPG:
P1000155.JPG

Comments:

  • Updated. Cables labeled following the ECR(EDMS 804188) -- Main.IvanMoyaMartinez - 09 Jan 2007 - 11:00

Edit, attach file or add comment


Created by Main.RobertoSaban on 2006-12-20 - 09:52

Incident Summary Report - Short on the MB circuit during TP4B

During the Electrical Quality Assurance (TP4B) of sector 78 a short to ground was identified. The short chronological report of the events is attached.

PDF short_21R7.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/short_21R7.pdf

PowerPoint FSavary.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/FSavary.ppt
Report to MARIC of the repair sequence
PowerPoint DBozzini.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x20x09x48/DBozzini.ppt
Report to MARIC of the ELQA tests

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-19 - 12:00

Preparation Cool-Down Sector 78-RAT 19th Dec

PDF RAT_20061219.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x19x11x13/RAT_20061219.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 15.12.06 DFBXG non-conformity on the current lead heater cables (P.Chambouvet-J-C.Guillaume)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
13 kA EE Systems
  • 16.12.06 Calibration of the Arc Detectors on the 13 kA EE System (K.Dahlerup-Petersen)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

MircoCoccoli - 19 Dec 2006 - 11:56

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-18 - 14:27

RAT for Point 5, Dec 18

PDF RAT_20061218.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x18x14x23/RAT_20061218.pdf

Comments: NEXT MEETING WILL BE HELD AT POINT 6 ON THURSDAY, 21

Open Issues for RR57/UJ56

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Open Issues for UA67/UJ67

  • 18.12 ERD rack powering
  • 18.12 EOD rack powering
  • 18.12 Connection of cable extremities on DQS
  • 18.12 HV test with circulating water
  • 18.12 Water-cooled cables supports
  • 18.12 PC powering
  • 18.12 Water circulation
  • 18.12 Elettas calibration
  • 18.12 Ventilation station
  • 18.12 CYNET01 and CYNET02 powering
  • 18.12 WorldFip cabling for PO
  • 18.12 Cooling station test by Roese
  • 18.12 Electrical protection IP2 on LEMs
  • 18.12 Cleaning of the areas

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-18 - 12:28

Preparation Cool-Down Sector 78-RAT 18thDec

PDF RAT_20061218.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x18x12x27/RAT_20061218.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-15 - 12:15

Preparation Cool-Down Sector 78-RAT 15th Nov

PDF RAT_20061215.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x15x12x14/RAT_20061215.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-12-14 - 18:25

Preparation Cool-Down Sector 78-RAT 14th Nov

PDF RAT_20061214.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x14x18x20/RAT_20061214.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-14 - 15:15

60A converter tests - Thu. 14th Dec.

Advancement of today:

  • connection of the half arc L8 finished: ready for ELQA.doc
  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for co activities.

Tomorrow activities:

  • TS-EL, still on resolution of the problem on the short circuit insulation in R7, no problems for co activities.
  • If it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for co activities.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-13 - 20:28

60A converter tests - Wed. 13th Dec.

JPEG P1000122.JPG:
P1000122.JPG

Advancement of today:

  • Pre-connection of the cable of the current leads: on C28-C32_L8
  • All 60A cables disconnected PC side:
  • Problem on the short circuit insulation on the 60A cables in R7 (see pic)

Tomorrow activities:

  • TS-EL, resolution of the problem on the short circuit insulation in R7, no problems for coactivities.
  • if it is possible: El_QA.doc 120A circuits in RR77 and UA83, no problems for coactivities.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-13 - 17:25

RAT Preparation Cool Down sec78 - Dec 13

PDF RAT_20061213.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x13x17x22/RAT_20061213.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 12.12.06 Short-to-ground in QBBI.A21R7 (Lyra side) (D.Bozzini, F.Seyvet)
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 12.12.06 Repositioning of 1 DC cable on DFBAO and of 2 DV cables on DFBAN (all 13kA) (J.C.Guillaume)
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 13 Dec 2006 - 17:22

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-12 - 20:28

60A converter tests - Tuesday 12th Dec.

Advancement Tue. 12 Dec.:

  • AB-PO activities ended in the whole arc
  • Polarity tests ended in the whole arc

Activities Wed. 13th Dec.:

  • TS-EL pre-conneciton in C32-34_L8 / C 12-34_R7; no problems for coactivities
  • if it is possible: El_QA.doc in C18-28 _L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-12 - 15:21

RAT for Point 5 in CCC, Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/RAT_20061212.pdf

PDF UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x12x15x08/UJ56-RR57_IR-Report_2006-12-12_24H-Burnin.pdf

Comments: NEXT MEETING WILL BE HELD ON DEC,18 AT POINT 5

Please, find also included the IR analysis rapport, kindly provided by Jeff Thomsen.

Open Issues

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 01.12 Squeeze test DONE (all the power converters involved have been tested)
  • 06.12 8-h heat run DONE
  • 07.12 Fire detection test DONE
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00 DONE
  • 08.12 Afternoon: Polarity Test_1st part TO BE RE-SCHEDULED
  • 11.12 10am: 24-h heat run DONE
  • 13.12 9.00: UPS test and Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part TO BE RE-SCHEDULED
  • 14.12 13kA energy extraction test
  • 15.12 14.00: New sequencer test

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-12 - 12:40

Preparation for Cool Down RAT sec78 - Dec 12

PDF RAT_20061212.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x12x12x38/RAT_20061212.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum
  • 12.12.06 One Pressure Gauge in DFBs cannot be read (F.Millet)
Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 08.12.06 DFBXG 100V test on current leads (S.Feher)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 12 Dec 2006 - 12:38

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-11 - 16:08

Preparation for Cool Down RAT - Dec 11

PDF RAT_20061211.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x11x16x04/RAT_20061211.pdf

Comments:

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 11 Dec 2006 - 16:07

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-11 - 13:18

RAT for Point5 in CCC, Dec 11

PDF RAT_20061211.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x11x13x11/RAT_20061211.pdf

Comments: Persons on call for assistance during the 24h heat run:

  • Hardware commissioning coordination: Mirko Pojer 162482
  • Power converters :
    • Responsable ABPO: David Nisbet 164223
    • (en cas de absence) Hugues Thiesen 164205
    • Numéro général ABPO-LHC-SCT-24H: 162333
    • Permanences ABPO: Journée 8h a 18 heures 11-12-2006 : Emmanuel Garde (16 4035) / Philippe Marchon Journée 18h a 20 heures 11-12-2006 : David Nisbet (164223) / Jeff Thomsen (165008) Nuit 20h 11-12-2006 a 2h00 12-12-2006 : Okbi Zenasni tel (16 3678) Nuit 2h a 8h 12-12-2006 : Cédric Caffier tel (16 4133) Journée 8h a 17 h 12-12-2006 : Emmanuel Garde (16 4035) / Philippe Marchon
  • Energy extraction system: Knud Dahlerup-Petersen 163279
  • DC cables: J. Kis (10h-18h), B. Lebeau (18h-00h), S. Cotte (00h-7h), J. Kis (7h-end of test) 164879

Open Issues

  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 01.12 Squeeze test DONE (all the power converters involved have been tested)
  • 06.12 8-h heat run DONE
  • 07.12 Fire detection test DONE
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00 DONE
  • 08.12 Afternoon: Polarity Test_1st part TO BE RE-SCHEDULED
  • 11.12 10am: 24-h heat run
  • 13.12 UPS test and Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part TO BE RE-SCHEDULED
  • 14.12 13kA energy extraction test.
  • 15.12 New sequencer test

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-08 - 17:54

60A converter tests - Friday 7th Dec.

Advancement of today:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • AB-PO main activity ended today;
  • Polarity tests on cells C24-26_R7

Monday activities:

  • AB-PO - tests on faulty cells: - C34_L8 - 1 converter in C32_R7 - 1 converter in C12_R7 - no problems for coactivities.
  • Polarity and conneciton in C28-32_L8 / C30-34_R7; no problems for coactivities
  • ELQA in C18-26_L8; co activities limited - a trolley is used

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-08 - 10:27

Preparation for Cool Down RAT - Dec 8

PDF RAT_20061208.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x08x10x17/RAT_20061208.pdf

IMPORTANT: PLEASE COMMENT ON THE OPEN ISSUES IF YOU FIND THEM INCORRECT, INCOMPLETE OR NOT-PERFECTLY UPDATED.

AC non-conformity

  • 07.12.06 Non conformity in DFBMH, DFBAN and DFBXG current lead heaters (P.Chambouvet-J-C.Guillaume)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling

  • 30.11.06 Non conformity temperature sensor cables need repair LSS7R (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBXG (P.Chambouvet)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Vacuum -

Cryogenics

  • 22.11.06 Cryo-valves remote control in CCC by AB/CO (F.Millet)
  • 22.11.06 Control of quench valve (F.Millet)
DFB Commissioning
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk50 (J-C.Guillaume & R.Jones)

MircoCoccoli - 08 Dec 2006 - 10:17

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-07 - 19:02

60A converter tests - Thursday 7th Dec.

Comments:

Advancement of today:

  • AB-PO activities on cells C12_R7 - C16-22_R7
  • Polarity tests on 120A converter alimenting cells C6-7-8-9-10_R7
  • Disconnection of the circuits of the 120A converter alimenting cells C6-7-8-9-10_R7
  • MME activities postponed to Monday

Tomorrow activities:

  • AB-PO activities on cells C24-30_R7
  • 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
  • ELQA tests - whole day / no co-activities

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-07 - 12:37

Preparation for Cool Down sec78 - Dec 7

PDF RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x07x12x29/RAT_20061207.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 4 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

MircoCoccoli - 07 Dec 2006 - 12:36

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-07 - 12:36

RAT for Point 5 in CCC, Dec 7

PDF RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x07x12x32/RAT_20061207.pdf

Comments: Next meeting in CCC on Monday Dec,11.

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
  • 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
  • 07.12 Cables re-distribution for 600 A
  • 07.12 Access key to electrical distribution room in UJ56

Heat Runs and Test Schedule

  • 07.12 Fire detection test
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00
  • 08.12 Afternoon: Polarity Test_1st part
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 13.12 Polarity Test_2nd part
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2006-12-06 - 20:51

60A converter tests - Wed. 6th Dec.

Advancement of today:

  • AB-PO performed its tests on C34_L8 - C34-32_R7 - C12-14_R7 Problems on converters on C34_L8 and C12_R7 - location to be re-visited
  • AB-PO/Polarity tests on 120A in RR77; circuits checked: RCBCH6, RCBCH7, RCBCH8, RCBCH9, RCBCH10 RCBCV6, RCBCV7, RCBCV8, RCBCV9, RCBCV10; Problems on the converters of the circuits RCBCH6 ,RCBCV6, RCBCV7 and RCBCH10, location to be re-visited
  • Polarity tests on C26-28_L8
  • TS-MME activities on C18-20-22-24-26_L8

Tomorrow activities:

  • AB-PO activities on cells C34_L8 - C12_R7 - C16-20_R7 no problems for co activities.
  • Polarity tests on C6-7-10_R7 no problems for co activities.
  • TS-MME activities on C6-7-8-9-10_R7 co activities may be problematic with BLM/BPM

Information:

  • 4 hr run of a set of converters foreseen for Friday

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-06 - 15:17

Preparation for Cool Down sec78 - Dec 6

PDF RAT_20061206_wslides_wList.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x06x15x14/RAT_20061206_wslides_wList.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)

Main.MircoCoccoli - 06 Dec 2006 - 15:16

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-06 - 12:47

RAT for Point 5 in CCC, Dec 6

PDF RAT_20061206.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x06x12x42/RAT_20061206.pdf

Comments: 8-H HEAT RUN ON GOING

Next meeting: Thursday, 7 at 8.30 in CCC.

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)

Heat Runs and Test Schedule

  • 07.12 Fire detection test
  • 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 13.12 Polarity Test
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-12-05 - 18:25

60A converter tests - Tuesday 5th Dec.

60A tests advancements

Advancement of today:

  • AB-PO performed its tests on C24-26-28-30-32_L8
  • Polarity tests on C22-24_L8
  • TS-MME activities on C12-14_L8
  • AT-MEL activities on cells C12-14L8
  • Positioning of the Cable in sector 7-8 finished at 11:30.

Tomorrow activities:

  • AB-PO activities on cells C34_L8 - C34-32-30-28_R7 no problems for co activities.
  • TS-MME activities on C18-20-22-24-26_L8 no problems for co activities
  • Polarity tests on C28-34_L8 no problems for co activities.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-05 - 13:50

Preparation for Cool Down sec78 - 5th Dec 06

PDF RAT_20061205_wslides.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x05x13x47/RAT_20061205_wslides.pdf
Minutes with Presentation by D.Bozzini attached

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 05 Dec 2006 - 13:47

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-05 - 12:58

Special RAT on SCT in RR57/UJ56 and ECR application, Dec 5

PDF RAT_20061205.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RAT_20061205.pdf

PowerPoint RQ4-RQ10-ECR.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RQ4-RQ10-ECR.ppt

Microsoft Excel Spreadsheet RR57-UJ56-8H-24H-PC_LIST.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x05x12x51/RR57-UJ56-8H-24H-PC_LIST.xls

Comments:

Please find enclosed the list of PC which will be powered during the 8h heat run and the 24h heat run. The list of steps to be implemented following the ERC (EDMS No.804188) is also included

Open Issues

  • 22.11 Fire detection – awaiting certification
  • 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-04 - 19:03

Preparation for Cool Down sec78 - RAT 4th Dec

PDF RAT_20061204.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/RAT_20061204.pdf

PDF Slide_w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x04x17x39/Slide_w49.pdf
Week 49 Usage Table

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 04.12.06 Access & Safety Cable Installation (full sector) (J-C.Guillaume; end by Dec 7th)
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (D.Bozzini, A.Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 04 Dec 2006 - 17:40

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-03 - 23:20

Sector 78 - Week 49 - Zones Usage

This tentative planning will be discussed at Monday 4th RAT meeting in Point 8.

Thanks, Mirco

PDF Plan78w49.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x03x23x13/Plan78w49.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-12-01 - 13:16

RAT - Preparation for Cool Down sec 78 - 1 Dec 06

PDF RAT_20061201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x01x13x13/RAT_20061201.pdf

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (A. Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 01 Dec 2006 - 13:13

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-12-01 - 11:49

RAT Point 5, Dec 1

PDF RAT_20061201.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x01x11x42/RAT_20061201.pdf

Comments: Next meeting oin Wednesday, Dec 6, at 8.30 in CCC

Open Issues

  • 23.11 Ripples in RQT12R5B1 power converter.
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Heat Runs and Test Schedule

  • 01.12 Squeeze test
  • 06.12 8-h heat run
  • 07.12 Polarity Test
  • 08.12 Full power ventilation tests
  • 11.12 10am: 24-h heat run
  • 12.12 UPS test.
  • 13.12 Electric cut (AUG simulation).
  • 14.12 13kA energy extraction test.

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-12-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login